### v2

# YITP-25-45

YITP-25-45
Family Unification in $SO(16)$ Grand Unification
Cheng-Wei Chiang1,2,3222chengwei@phys.ntu.edu.tw, Tianjun Li4,5,6333tli@itp.ac.cn, and Naoki Yamatsu7111naoki.yamatsu@yukawa.kyoto-u.ac.jp
1Department of Physics and Center for Theoretical Physics, National Taiwan University, Taipei, Taiwan 10617, R.O.C.
2Physics Division, National Center for Theoretical Sciences, Taipei, Taiwan 10617, R.O.C.
3Kavli IPMU (WPI), UTIAS, The University of Tokyo, Kashiwa, 277-8583, Japan
4CAS Key Laboratory of Theoretical Physics, Institute of Theoretical Physics,
Chinese Academy of Sciences, Beijing 100190, P. R. China
5School of Physical Sciences, University of Chinese Academy of Sciences,
No. 19A Yuquan Road, Beijing 100049, P. R. China
6School of Physics, Henan Normal University, Xinxiang 453007, P. R. China
7Yukawa Institute for Theoretical Physics, Kyoto University,
Kitashirakawa Oiwakecho, Sakyo-ku, Kyoto 606-8502, Japan
1 Introduction
The existence of three chiral generations of quarks and leptons is one of the most mysterious facts in particle physics. In addition, the mass hierarchy among different flavors of fermions generated by the Higgs mechanism via their corresponding Yukawa couplings strongly calls for a better understanding of the underlying structure of Nature. There are many attempts to understand the origin of chiral generations and/or their mass hierarchy by considering, e.g., the so-called horizontal symmetry (or family symmetry) in four-dimensional (4D) theories [1, 2, 3, 4, 5, 6, 7, 8, 9], as well as employing special geometrical structures in higher-dimensional theories [10, 11].
It is well-known that the quarks and leptons for each generation in the Standard Model (SM) can be unified into one multiplet (or two multiplets) in grand unified theories (GUTs) in the 4D [12, 13, 14, 15, 16, 17] and higher-dimensional frameworks [18, 19, 20, 21, 22, 23, 24, 25, 26].
There are some attempts to unify GUT and family groups into a larger GUT group in 4D and higher-dimensional theories [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. In particular, models based on $SU(19)$ GUT gauge symmetry are discussed in the 4D framework [32, 37] and the six-dimensional (6D) framework by one of the authors [36]. In these models, the three chiral generations of the SM fermions are unified into a fermion multiplet in the second-rank antisymmetric tensor representation ${\bf 171}$ of $SU(19)$. In such models, $SU(19)$ is broken to a maximal regular subgroup $SU(16)\times SU(3)\times U(1)$, and $SU(16)$ is further broken to a maximal special subgroup $SO(10)$. The branching rule of ${\bf 171}$ under $SU(19)\supset SO(10)\times SU(3)\times U(1)$ is given by [39]
${\bf 171}=({\bf 16,3})(-13)\oplus{\bf(120,1)}(6)\oplus{(\bf 1,\overline{3})}(-32),$ (1.1)
where the $SO(10)$ ${\bf 120}$ representation is real. That is, only the first term is a complex representation of $SO(10)$. Therefore, when we introduce a fermion in ${\bf 171}$ of $SU(19)$, only the three chiral generations of SM fermions appear after we take into account $SU(3)$ and $U(1)$ breaking effects. For more information about this type of symmetry breaking, see Refs. [39, 40, 41, 36].
For $SO(N)$ GUT gauge symmetry, the models are inconsistent with low-energy phenomena because additional generations or mirror particles would appear in the 4D framework. There have been discussions aiming at unifying the three chiral generations of SM fermions in the framework of 5-dimensional (5D) orbifold spacetime [35]. In Ref. [35], unification of the SM gauge symmetry and family symmetry based on $SO(18)$ GUT gauge symmetry is discussed in the 5D spacetime of $M^{4}\times S_{1}/(\mathbb{Z}_{2}\times\mathbb{Z}_{2}^{\prime})$. The $SO(18)$ gauge symmetry is broken to $SO(10)\times SO(8)$ by the orbifold. Besides, the confinement of $SO(5)$, which is a subgroup of $SO(8)$, is used along with an orbifold space structure to remove the mirror SM fermions. One of the most subtle points in this model is the use of the branching rule for the vector representation as the branching rule for the spinor representation of $SO(8)$. Furthermore, the $SO(18)$ gauge symmetry is used in this model for this reason to achieve the $SO(5)$ confinement. If we do not need to use confinement, the $SO(16)$ gauge symmetry should be sufficient to unify the three chiral generations of SM fermions. This is because the branching rules of $SO(16)\supset SO(10)\times SU(3)\times U(1)$ for spinor representations ${\bf 128}$ and ${\bf 128^{\prime}}$ are given by
${\bf 128}=({\bf 16,3})(1)\oplus({\bf 16,1})(-3)\oplus\oplus({\bf\overline{16},\overline{3}})(-1)\oplus({\bf\overline{16},1})(3),$
${\bf 128^{\prime}}=({\bf 16,\overline{3}})(-1)\oplus({\bf 16,1})(3)\oplus({\bf\overline{16},3})(1)\oplus({\bf\overline{16},1})(-3).$ (1.2)
(For more information about Lie algebras and their subalgebras, see, e.g., Ref. [39].)
In this paper, we will propose a new model based on $SO(16)$ GUT gauge symmetry that unifies the three chiral generations of SM fermions into a 6D Weyl fermion in a spinor representation of $SO(16)$ without exotic fermions on the 6D spacetime $M^{4}\times D^{2}/\mathbb{Z}_{N}$ $(N\geq 9)$, where $D^{2}$ is a 2-dimensional (2D) disk and there is a fixed point at the origin [42]. The 6D $SO(16)$ gauge anomaly is canceled by the vectorlike nature of the model, and the 4D gauge anomalies are canceled by introducing suitable 4D localized fermions at the fixed point. The gauge coupling constant of $SO(16)$ is asymptotically free.
This paper is organized as follows. In Sec. 2, we introduce the 6D spacetime $M^{4}\times D^{2}/\mathbb{Z}_{N}$ and $SO(16)$ gauge symmetry along with its breaking. In Sec. 3, we discuss the relations among the symmetries, gauge bosons, and orbifolding the extra space $D^{2}/\mathbb{Z}_{N}$. In Sec. 4, we show how to realize the three chiral generations of SM fermions without exotic fermions on $M^{4}\times D^{2}/\mathbb{Z}_{N}$. In Sec. 5, we examine how to realize 6D and 4D gauge anomaly cancellation in the model. In Sec. 6, we demonstrate how to realize the breakdown of $SO(16)$ GUT symmetry to the SM gauge symmetry. In Sec. 7, we analyze the renormalization group equation (RGE) for the gauge coupling constant of $SO(16)$. Section 8 summarizes and discusses our findings in this work.
2 6D spacetime and $SO(16)$ gauge symmetry
Let’s consider the 6D spacetime $M^{4}\times D^{2}/\mathbb{Z}_{N}$ with Cartesian coordinates $x^{\mu}$ ($\mu=0,1,2,3$) and circular coordinates $(r,\theta)$ or a complex coordinate $z=re^{i\theta}$, where $D^{2}$ is a 2-dimensional (2D) disk and there is a fixed point at the origin $r=0$ [42]. We take the radius of the disc $D^{2}$ as $R$. A $\mathbb{Z}_{N}$ symmetry on $D^{2}$ is defined as
$\omega=e^{i\frac{2\pi}{N}},$ (2.1)
where the generator for $\mathbb{Z}_{N}$, $\Omega$, satisfies $\Omega^{N}=1$. The metric of the 6D spacetime is given by
$ds^{2}=\eta_{\mu\nu}dx^{\mu}dx^{\nu}+dr^{2}+r^{2}d\theta^{2},$ (2.2)
where $\eta_{\mu\nu}=\mbox{diag}(-1,1,1,1)$.
In the following, we discuss the gauge fields associated with the $SO(16)$ generators $T^{\alpha}$ ($\alpha=1,2,...,120$), which are expressed using the Pauli matrices $\sigma_{a}$ ($a=1,2,3$) and $8\times 8$ symmetric and anti-symmetric matrices $S_{8}$ and $A_{8}$ as [43]:
$T^{\alpha}=\{\sigma_{0},\sigma_{1},\sigma_{3}\}\otimes A_{8},\ \ \sigma_{2}\otimes S_{8},$ (2.3)
where $\sigma_{0}$ is the $2\times 2$ identity matrix. The numbers of $S_{8}$ and $A_{8}$ matrices are $36$ and $28$, respectively. (The total number of the generators is therefore $3\times 28+36=120$.) Denoting the corresponding components of the gauge field as $A_{M}^{\sigma_{x}}$ $(x=0,1,2,3)$, we can express the gauge field $A_{M}$ as
$A_{M}=\left(\begin{array}[]{cc}A_{M}^{\sigma_{0}}+iA_{M}^{\sigma_{3}}&A_{M}^{\sigma_{1}}+iA_{M}^{\sigma_{2}}\\ A_{M}^{\sigma_{1}}-iA_{M}^{\sigma_{2}}&A_{M}^{\sigma_{0}}-iA_{M}^{\sigma_{3}}\\ \end{array}\right).$ (2.6)
Here we consider $SO(16)$ symmetry breaking by the $\mathbb{Z}_{N}$ orbifold. First, the unbroken generators $T^{a}$ and the broken generators $T^{\hat{a}}$ satisfy respectively
$R_{\Omega}T^{a}R_{\Omega}^{-1}=+T^{a},\ \ \ R_{\Omega}T^{\hat{a}}R_{\Omega}^{-1}\not=T^{\hat{a}},$ (2.7)
where $R_{\Omega}$ denotes an element in the adjoint representation of the bulk gauge group $G$ that satisfies $R_{\Omega}^{N}=I_{n}$, where $I_{n}$ is the $n\times n$ identity matrix. For a generic bulk multiplet $\Phi$ in a representation of $G$, we have
$\Omega\Phi(x^{\mu},z,\bar{z})=\Phi(x^{\mu},\omega z,\bar{\omega}\bar{z})=\eta_{\Phi}\left(R_{\Omega}\right)^{\ell_{\Phi}}\Phi(x^{\mu},z,\bar{z})\left(R_{\Omega}^{-1}\right)^{m_{\Phi}},$ (2.8)
where $\bar{\omega}:=\omega^{-1}$, $\eta_{\Phi}\subset\mathbb{Z}_{N}$, and $\ell_{\Omega}$ and $m_{\Omega}$ are non-negative integers depending on the representation of $G$. The boundary condition at $r=R$ in $D^{2}/\mathbb{Z}_{N}$ can be chosen as the Dirichlet or Neumann boundary condition. However, if the Dirichlet boundary condition is chosen, there will be no zero mode, so in the ensuing discussion, the Neumann boundary condition will be chosen to allow the existence of zero modes. (For details, see, e.g., Refs. [28, 42].) We will apply this discussion for the $SO(16)$ gauge field $A_{\mu}$ in the 4D spacetime below, noting that different choices of $R_{\Omega}$ will lead to different symmetry-breaking schemes.
First, we consider that $SO(16)$ is broken to $SU(8)\times U(1)_{a}$ by the $\mathbb{Z}_{N}$ orbifold. The branching rules of $SO(16)\supset SU(8)\times U(1)_{a}$ for some representations are given by [39]
${\bf 16}$ $=({\bf 8})(1)\oplus({\bf\overline{8}})(-1),$
${\bf 120}$ $=({\bf 63})(0)\oplus({\bf 1})(0)\oplus({\bf 28})(2)\oplus({\bf\overline{28}})(-2),$
${\bf 128}$ $=({\bf 56})(-1)\oplus({\bf\overline{56}})(1)\oplus({\bf 8})(-3)\oplus({\bf\overline{8}})(3),$
${\bf 128^{\prime}}$ $=({\bf 70})(0)\oplus({\bf 28})(-2)\oplus({\bf\overline{28}})(2)\oplus({\bf 1})(4)\oplus({\bf 1})(-4).$ (2.9)
This symmetry-breaking pattern is realized by taking the following matrix:
$R_{\Omega}=\mbox{exp}\left[i\frac{2\pi}{N}\sigma_{2}\otimes(-a)I_{8}\right],$ (2.10)
where $a=1,2,\cdots,N-1$ (mod. $N$), and the exponents are proportional to the $U(1)_{a}$ charges. In this case, the gauge field $A_{\mu}$ is transformed under $\mathbb{Z}_{N}$ as
$A_{\mu}=\left(\begin{array}[]{cc}+A_{\mu}^{\sigma_{0}}+iA_{\mu}^{\sigma_{3}}&+A_{\mu}^{\sigma_{1}}+iA_{\mu}^{\sigma_{2}}\\ +A_{\mu}^{\sigma_{1}}-iA_{\mu}^{\sigma_{2}}&+A_{\mu}^{\sigma_{0}}-iA_{\mu}^{\sigma_{3}}\\ \end{array}\right)\ \to\ R_{\Omega}A_{\mu}R_{\Omega}^{-1}=\left(\begin{array}[]{cc}+A_{\mu}^{\sigma_{0}}-iA_{\mu}^{\sigma_{3}}&-A_{\mu}^{\sigma_{1}}+iA_{\mu}^{\sigma_{2}}\\ -A_{\mu}^{\sigma_{1}}-iA_{\mu}^{\sigma_{2}}&+A_{\mu}^{\sigma_{0}}+iA_{\mu}^{\sigma_{3}}\\ \end{array}\right).$ (2.15)
The $\sigma_{0}$ and $\sigma_{2}$ component fields $A_{\mu}^{\sigma_{0}}$ and $A_{\mu}^{\sigma_{2}}$ are invariant under $\mathbb{Z}_{N}$, while the $\sigma_{1}$ and $\sigma_{3}$ component fields $A_{\mu}^{\sigma_{1}}$ and $A_{\mu}^{\sigma_{3}}$ are not invariant. Therefore, $SO(16)$ is broken to $SU(8)\times U(1)_{a}$.
Next, we consider $SO(16)$ breaking to $SU(5)\times SU(3)\times U(1)_{a}\times U(1)_{b}$ by the $\mathbb{Z}_{N}$ orbifold. The branching rules of $SU(8)\supset SU(5)\times SU(3)\times U(1)_{b}$ for some representations are given by [39]
${\bf 8}$ $=({\bf 5,1})(3)\oplus({\bf 1,3})(-5),$
${\bf\overline{8}}$ $=({\bf\overline{5},1})(-3)\oplus({\bf 1,\overline{3}})(5),$
${\bf 28}$ $=({\bf 10,1})(6)\oplus({\bf 5,3})(-2)\oplus({\bf 1,\overline{3}})(-10),$
${\bf\overline{28}}$ $=({\bf\overline{10},1})(-6)\oplus({\bf\overline{5},\overline{3}})(2)\oplus({\bf 1,3})(10),$
${\bf 56}$ $=({\bf\overline{10},1})(9)\oplus({\bf 10,3})(1)\oplus({\bf 5,\overline{3}})(-7)\oplus({\bf 1,1})(-15),$
${\bf\overline{56}}$ $=({\bf 10,1})(-9)\oplus({\bf\overline{10},\overline{3}})(-1)\oplus({\bf\overline{5},3})(7)\oplus({\bf 1,1})(15),$
${\bf 63}$ $=({\bf 24,1})(0)\oplus({\bf 1,8})(0)\oplus({\bf 1,1})(0)\oplus({\bf 5,\overline{3}})(8)\oplus({\bf\overline{5},3})(-8),$
${\bf 70}$ $=({\bf 10,\overline{3}})(-4)\oplus({\bf\overline{10},3})(4)\oplus({\bf 5,1})(-12)\oplus({\bf\overline{5},1})(12).$ (2.16)
We find that the following matrix realizes $SO(16)$ breaking to $SU(5)\times SU(3)\times U(1)_{a}\times U(1)_{b}$:
$R_{\Omega}=\mbox{exp}\left[i\frac{2\pi}{N}\sigma_{2}\otimes\left\{\left(-a-3b\right)I_{5}\oplus\left(-a+5b\right)I_{3}\right\}\right],$ (2.17)
where $|a+3b|,|a-5b|\in\mathbb{Z}_{\not=0}$, and the exponents are proportional to the $U(1)_{a}$ and $U(1)_{b}$ charges.
Finally, we consider $SO(16)$ breaking to $SO(10)\times SU(3)\times U(1)$ by the $\mathbb{Z}_{N}$ orbifold. The branching rules of $SO(16)\supset SO(10)\times SU(3)\times U(1)$ for some representations are already given in Eq. (1.2) and by [39]
${\bf 16}$ $=({\bf 10,1})(0)\oplus({\bf 1,\overline{3}})(2)\oplus({\bf 1,3})(-2),$
${\bf 120}$ $=({\bf 45,1})(0)\oplus({\bf 1,8})(0)\oplus({\bf 1,1})(0)\oplus({\bf 1,3})(4)\oplus({\bf 1,\overline{3}})(-4)\oplus({\bf 10,3})(-2)\oplus({\bf 10,\overline{3}})(2).$ (2.18)
This symmetry-breaking pattern is realized by choosing $a=-3\ell/4$ and $b=\ell/4$ $(\ell\in\mathbb{Z}_{\not=0})$ in Eq. (2.17) as
$R_{\Omega}=\mbox{exp}\left[i\frac{2\pi}{N}\sigma_{2}\otimes\left\{0\cdot I_{5}\oplus 2\ell\cdot I_{3}\right\}\right],$ (2.19)
where the exponents are proportional to the $U(1)$ charges. This matrix breaks $SO(16)$ to $SO(10)\times SU(3)\times U(1)$. Note that if we choose special values of $N$ and $\ell$, then $SO(16)$ is unbroken for $\ell=0$ (mod. $N/2$), and $SO(16)$ is broken to $SO(10)\times SO(6)\simeq SO(10)\times SU(4)$ for $\ell=N/4$ (mod. $N/2$), where $\ell\in\mathbb{Z}_{\not=0}$. We will consider the $SO(16)$ breaking to $SO(10)\times SU(3)\times U(1)$ in the following sections.
3 Gauge symmetry and gauge boson
In this section, we discuss the conditions that $SO(16)$ GUT gauge symmetry can be broken to $SO(10)\times SU(3)\times U(1)$. We also find the conditions that no extra scalar particles emerge from the extra-dimensional components of the gauge boson.
The gauge boson transforms under $\Omega$ as [28, 42]
$\Omega A_{\mu}(x^{\mu},z,\bar{z})=A_{\mu}(x^{\mu}\omega z,\bar{\omega}\bar{z})=R_{\Omega}A_{\mu}(x^{\mu},z,\bar{z})R_{\Omega}^{-1},$
$\Omega A_{z}(x^{\mu},z,\bar{z})=A_{z}(x^{\mu},\omega z,\bar{\omega}\bar{z})=\omega^{-1}R_{\Omega}A_{z}(x^{\mu},z,\bar{z})R_{\Omega}^{-1}.$ (3.1)
The phase factors of the components of the gauge field $A_{M}$ $(M=\mu,z)$ are summarized in Table 1, where $G_{10,3,1}:=SO(10)\times SU(3)\times U(1)$.
$G_{10,3,1}$ $A_{\mu}$ $A_{z}$
$({\bf 45,1})(0)$ $1$ $\omega^{-1}$
$({\bf 1,8})(0)$ $1$ $\omega^{-1}$
$({\bf 1,1})(0)$ $1$ $\omega^{-1}$
$({\bf 1,3})(+4)$ $\omega^{+4\ell}$ $\omega^{-1+4\ell}$
$({\bf 1,\overline{3}})(-4)$ $\omega^{-4\ell}$ $\omega^{-1-4\ell}$
$({\bf 10,\overline{3}})(+2)$ $\omega^{+2\ell}$ $\omega^{-1+2\ell}$
$({\bf 10,{3}})(-2)$ $\omega^{-2\ell}$ $\omega^{-1-2\ell}$
Table 1: Phase factors of the components of the gauge field $A_{M}$ $(M=\mu,z)$, where $\ell$ is an integer.
From Table 1, we find that there are the following symmetry breaking patterns, depending on the value of $\ell$:
$SO(16)\to\left\{\begin{array}[]{ll}SO(16)&\mbox{for}\ \ell=0\ (\mbox{mod.}\ N/2),\\ SO(10)\times SU(4)&\mbox{for}\ \ell=N/4\ (\mbox{mod.}\ N/2),\\ SO(10)\times SU(3)\times U(1)&\mbox{for the others}.\end{array}\right.$ (3.5)
We focus on the symmetry-breaking pattern $SO(16)\to SO(10)\times SU(3)\times U(1)$ in Eq. (3.5), where the three chiral generations of SM fermions can be unified into one fermion multiplet. In this case, we can take $\ell\not=0$ (mod. $N/4$). Here, we further impose the condition that no extra zero mode appears from the extra-dimensional components of the gauge field. From Table 1, we find the following constraints for $\ell$ and $N$ to avoid the unwanted zero modes:
±$\pm 2\ell,\ \pm 4\ell,\ -1\pm 2\ell,\ -1\pm 4\ell\not=0\ (\mbox{mod.}\ N).$ (3.6)
Obviously, we cannot take $\ell=0$. We next consider the $\ell=1$ case. From Eq. (3.6), we find
[…]
This leads to the requirement that $N\geq 6$. For $\ell=2$, from Eq. (3.6), we find
[…]
This leads to $N=6$ or $N\geq 10$. We can also consider $\ell\geq 3$ and negative integers in the same manner.
4 Unification of the three chiral generations of SM fermions
We introduce 6D positive and negative Weyl fermions $\Psi_{\bf 128}^{\alpha=1}(x,z,\bar{z})$ and $\Psi_{\bf 128}^{\alpha=2}(x,z,\bar{z})$ in a spinor representation ${\bf 128}$ of $SO(16)$. The purpose of having the “positive” and “negative” Weyl fermions is to render a vectorlike fermion in the 6D gauge theory. These 6D Weyl fermions can be decomposed into 4D Weyl fermions as
$\Psi_{\bf 128}^{\alpha}(x,z,\bar{z})=\left(\begin{array}[]{c}\psi_{L{\bf 128}}^{\alpha}\\ \psi_{R{\bf 128}}^{\alpha}\\ \end{array}\right),$ (4.3)
where $\psi_{L{\bf 128}}^{\alpha}$ and $\psi_{R{\bf 128}}^{\alpha}$ stand for 4D left- and right-handed Weyl fermions, respectively. The transformations of fermions under $\Omega$ are given by [28, 42]
$\Omega\Psi_{\bf 128}^{\alpha}(x^{\mu},z,\bar{z})=\Psi_{\bf 128}^{\alpha}(x^{\mu},\omega z,\bar{\omega}\bar{z})=\eta_{\Psi^{\alpha}}R_{\Omega}^{\rm sp}\Psi_{\bf 128}^{\alpha}(x^{\mu},z,\bar{z}),$ (4.4)
[…]
$\Psi_{\bf 128}^{\alpha}$
$G_{10,3,1}$ Left Leftc
$({\bf 16,3})(+1)$ $\eta_{\Psi^{\alpha}}\omega^{+\ell}$ $\eta_{\Psi^{\alpha}}^{\prime}\omega^{+\ell}$
$({\bf 16,1})(-3)$ $\eta_{\Psi^{\alpha}}\omega^{-3\ell}$ $\eta_{\Psi^{\alpha}}^{\prime}\omega^{-3\ell}$
$({\bf\overline{16},\overline{3}})(-1)$ $\eta_{\Psi^{\alpha}}\omega^{-\ell}$ $\eta_{\Psi^{\alpha}}^{\prime}\omega^{-\ell}$
$({\bf\overline{16},1})(+3)$ $\eta_{\Psi^{\alpha}}\omega^{+3\ell}$ $\eta_{\Psi^{\alpha}}^{\prime}\omega^{+3\ell}$
Table 2: Phase factors of the components of the fermions $\Psi_{\bf 128}^{\alpha}$ $(\alpha=1,2)$, where $\eta_{\Psi^{\alpha}}\eta_{\Psi^{\alpha}}^{\prime}=\omega$.
Here we consider the case of $SO(16)$ breaking to $SO(10)\times SU(3)\times U(1)$, where we take $\ell\not=0$ (mod. $N/4$). For $\Psi_{\bf 128}^{\alpha=1}$ with $\eta_{\Psi^{1}}=\omega^{-\ell}$, we find the following constraints for $\ell$ from the left table in Table 3 to realize the three chiral generations of SM fermions without producing exotic fermions:
$-4\ell,\ \pm 2\ell,\ \pm 2\ell+1,\ 1,\ 4\ell+1\not=0\ (\mbox{mod.}\ N).$ (4.5)
Also, we need to take $\eta_{\Psi^{2}}(=\omega^{k})$ not to produce zero modes for $\Psi_{\bf 128}^{\alpha=2}$. From the right table in Table 3, we find the following constraints for $k$ and $\ell$:
$k\pm\ell,\ k\pm 3\ell,\ -k\pm\ell+1,\ -k\pm 3\ell+1\not=0\ (\mbox{mod.}\ N).$ (4.6)
$\Psi_{\bf 128}^{\alpha=1}$ $\Psi_{\bf 128}^{\alpha=2}$
$G_{10,3,1}$ Left Leftc
$({\bf 16,3})(+1)$ $1$ $\omega^{+2\ell+1}$
$({\bf 16,1})(-3)$ $\omega^{-4\ell}$ $\omega^{-2\ell+1}$
$({\bf\overline{16},\overline{3}})(-1)$ $\omega^{-2\ell}$ $\omega^{+1}$
$({\bf\overline{16},1})(+3)$ $\omega^{+2\ell}$ $\omega^{+4\ell+1}$
$G_{10,3,1}$ Left Leftc
$({\bf 16,3})(+1)$ $\omega^{+k+\ell}$ $\omega^{-k+\ell+1}$
$({\bf 16,1})(-3)$ $\omega^{+k-3\ell}$ $\omega^{-k-3\ell+1}$
$({\bf\overline{16},\overline{3}})(-1)$ $\omega^{+k-\ell}$ $\omega^{-k-\ell+1}$
$({\bf\overline{16},1})(+3)$ $\omega^{+k+3\ell}$ $\omega^{-k+3\ell+1}$
Table 3: Phase factors of the components of the fermions $\Psi_{\bf 128}^{\alpha}$ $(\alpha=1,2)$, where $\eta_{\Psi^{\alpha}}\eta_{\Psi^{\alpha}}^{\prime}=\omega$. $\eta_{\Psi^{1}}=\omega^{-\ell}$ and $\eta_{\Psi^{2}}=\omega^{+k}$.
We need to find a set of $\ell$, $k$ and $N$ that simultaneously satisfy Eqs. (4.5) and (4.6). There turns out to be an infinite number of allowed sets of $\ell$, $k$ and $N$. Hence, we will only consider a few simple examples. First, we set $\ell=1$. In this case, Eq. (4.5) gives
[…]
We thus need to choose $N\geq 6$. On the other hand, Eq. (4.6) has
$k\pm 1,\ k\pm 3,\ -k+2,\ -k,\ -k+4,\ -k-2\not=0\ (\mbox{mod.}\ N).$ (4.8)
Taking the example of $k=-4$, we have the condition as
[…]
This leads to the requirement $N\geq 9$. In other words, we have now found a possible set of parameters, $(\ell,k,N)=(1,-4,9)$.
We next try $\ell=2$. In this case, Eq. (4.5) gives
[…]
This leads to $N=6,7$ or $N\geq 10$. On the other hand, Eq. (4.6) requires
$k\pm 2,\ k\pm 6,\ -k+3,\ -k-1,\ -k+7,\ -k-5\not=0\ (\mbox{mod.}\ N).$ (4.11)
Taking $k=0$ as an example, we have the requirement
[…]
This then leads to $N=4$ or $N\geq 8$. Combining the two conditions, we need to take $N\geq 10$. Thus, we have found another possible set of parameters: $(\ell,k,N)=(2,0,10)$.
In summary, to unify the three chiral generations of SM fermions without the occurrence of exotic fermions, one should take $(N,\eta_{\Psi^{1}},\eta_{\Psi^{2}})=(9,\omega^{-1},\omega^{-4})$, $(10,\omega^{-2},1)$, $\dots$ In such cases, only the fermions in the $({\bf 16,3})(1)$ representation of $SO(10)\times SU(3)\times U(1)$ have zero modes.
5 6D and 4D gauge anomaly cancellation
From the above discussions, it is clear that one choice of having the three chiral generations of SM fermions is to take $(N,\eta_{\Psi^{1}},\eta_{\Psi^{2}})=(9,\omega^{-1},\omega^{-4})$. In this case, we obtain the 4D left-handed fermions in $({\bf 16,3})(1)$ of $SO(10)\times SU(3)\times U(1)$ from a single 6D fermion $\Psi_{\bf 128}^{\alpha=1}$ without having exotic fermions. In the following, we discuss 6D and 4D gauge anomaly cancellation in this scenario.
First, we consider the 6D $SO(16)$ gauge anomaly cancellation. Our model includes 6D Weyl fermions and thus can in principle induce 6D gauge anomalies. By introducing one 6D positive Weyl fermion in 128 of $SO(16)$ and another 6D negative Weyl fermion in 128 of $SO(16)$, the fermions in the model are vectorlike in the 6D gauge theory. Therefore, the 6D $SO(16)$ gauge anomaly is trivially canceled.
Next, we discuss how to cancel the 4D gauge anomalies. As is well known, there are two types of 4D gauge anomalies: pure gauge anomalies and mixed gauge anomalies. Pure gauge anomalies are determined by the group structure in the case of semi-simple Lie groups, and there are no 4D pure gauge anomalies except for the $SU(n)$ and $U(1)$ groups [44, 45, 46, 39]. In other words, in the present case the pure gauge quantum anomaly is only in the $SU(3)$ and $U(1)$ symmetries, i.e., the $SU(3)\mbox{-}SU(3)\mbox{-}SU(3)$ and $U(1)\mbox{-}U(1)\mbox{-}U(1)$ gauge anomalies. There are only three mixed gauge anomalies, i.e., the $SO(10)\mbox{-}SO(10)\mbox{-}U(1)$, $SU(3)\mbox{-}SU(3)\mbox{-}U(1)$, and $\mbox{grav}\mbox{-}\mbox{grav}\mbox{-}U(1)$ anomalies. Furthermore, the pure gauge anomalies can be written in terms of third-order Dynkin indices or anomaly coefficients and the mixed gauge anomalies can be written in terms of second-order Dynkin indices and $U(1)$ charges, as follows:
$A_{\rm bulk}^{SU(3)\mbox{-}SU(3)\mbox{-}SU(3)}=16\times A_{SU(3)}{(\bf 3)}=+16,$
$A_{\rm bulk}^{U(1)\mbox{-}U(1)\mbox{-}U(1)}=16\times 3\times(+1)^{3}=+48,$
$A_{\rm bulk}^{SO(10)\mbox{-}SO(10)\mbox{-}U(1)}=3\times T_{SO(10)}{(\bf 16)}\times(+1)=+6,$
$A_{\rm bulk}^{SU(3)\mbox{-}SU(3)\mbox{-}U(1)}=16\times T_{SU(3)}{(\bf 3)}\times(+1)=+8,$
$A_{\rm bulk}^{\mbox{grav}\mbox{-}\mbox{grav}\mbox{-}U(1)}=16\times 3\times T_{SO(4)}{(\bf 2,1)}\times(+1)=+24,$ (5.1)
where $A_{SU(3)}{(\bf 3)}=+1$, $T_{SO(10)}{(\bf 16)}=2$, $T_{SU(3)}{(\bf 3)}=1/2$, and $T_{SO(4)}{({\bf 2,1})}=1/2$ [39].
We now examine what kind of 4D fields should be introduced to cancel out the gauge anomaly generated by the 6D fields given in Eq. (5.1). It turns out to be impossible to uniquely determine the required set of 4D fields solely by the requirement that the 4D gauge anomalies arising from the 6D fields cancel out. In the following, we provide an explicit example. One way is to introduce 4D localized left-handed fermions in $({\bf 10,\overline{3}})(-1)$ and $({\bf 1,3})(+1)$ of $SO(10)\times SU(3)\times U(1)$, denoted by $\psi_{L({\bf 10,\overline{3}})(-1)}^{(\beta)}$ $(\beta=1,2)$ and $\psi_{L({\bf 1,{3}})(+1)}^{(\gamma)}$ $(\gamma=1,2,3,4)$, respectively. In the case, the $SO(10)\mbox{-}SO(10)\mbox{-}U(1)$ mixed gauge anomaly is canceled by $\psi_{L({\bf 10,\overline{3}})(-1)}^{(\beta)}$; the remaining anomalies are canceled by the combination of $\psi_{L({\bf 10,\overline{3}})(-1)}^{(\beta)}$ and $\psi_{L({\bf 1,{3}})(+1)}^{(\gamma)}$. When we take into account the vacuum expectation values (VEVs) of the $SU(3)$- and $U(1)$-breaking scalar fields, these 4D fields obtain masses that do not appear at low energies and will not be discussed further in this paper.
6 Symmetry breaking via the VEVs of scalars
Next, we consider the scalar content to realize the SM gauge symmetry at low energies. The set of scalar fields, which properly break the $SO(16)$ symmetry into $G_{\rm SM}$, is also not uniquely determined. Here we provide a working example by introducing localized scalars $\phi_{{\bf(10,1)}(0)}$, $\phi_{{\bf(45,1})(0)}$, $\phi_{{\bf(16,1)}(0)}$, $\phi_{{\bf(1,3)}(0)}$, $\phi_{{\bf(1,6)}(+2)}$, and $\phi_{{\bf(1,1)}(+1)}$ to realize $G_{\rm SM}$ at low energies. In this scenario, the symmetry-breaking pattern is given as follows:
$\begin{array}[]{rll}SO(16)\ \to&SO(10)\times SU(3)\times U(1)&\mbox{by the non-trivial space structure}\\ \to&SO(10)&\mbox{by the VEVs}\ \langle\phi_{{\bf(1,3)}(0)}\rangle,\langle\phi_{{\bf(1,6)}(+2)}\rangle,\langle\phi_{{\bf(1,1)}(+1)}\rangle\not=0\\ \to&G_{\rm SM}&\mbox{by the VEVs}\ \langle\phi_{{\bf(45,1)}(0)}\rangle,\langle\phi_{{\bf(16,1)}(0)}\rangle\not=0\\ \to&SU(3)_{C}\times U(1)_{\rm EM}&\mbox{by the VEV}\ \langle\phi_{{\bf(10,1)}(0)}\rangle\not=0,\\ \end{array}$ (6.5)
where we assume that the Kaluza-Klein (KK) mass scales $1/R$ and the non-vanishing VEVs are at the order of $M_{\rm GUT}\simeq O(10^{16}-10^{17})$ GeV except the VEV of the SM Higgs $\langle\phi_{{\bf(10,1)}(0)}\rangle$. By assuming symmetry breaking with the scalar field VEVs given above, all the introduced 4D fermions will attain large mass and will not appear at low energies. The $SO(10)$ GUT symmetry breaking to the SM gauge symmetry has been thoroughly examined in, for example, Refs. [47, 48, 49, 50, 39] and will not be discussed further here.
7 Renormalization group equation for gauge coupling constant
Here we analyze the renormalization group equation (RGE) evolution of the $SO(16)$ gauge coupling constant at energy scales above the KK scale by using a KK-mode expanded 4D effective theory as discussed in Ref. [51]. Since the 4D fermions and scalar bosons do not have KK modes, they do not change the qualitative properties of the RGE evolution above the KK scale. In the following discussions, we will focus on the RGE evolution affected by the KK particles from the 6D fields.
We first summarize some basic results of the RGEs for 4D gauge coupling constants. The RGE for the $SO(16)$ gauge coupling constant $g_{SO(16)}$ is given in, e.g., Refs. [50, 39, 51]. The RGE for the $SO(16)$ fine structure constant $\alpha_{SO(16)}(\mu):=g_{SO(16)}^{2}(\mu)/4\pi$ at energy scale $\mu$ is given by
$\frac{d}{d\mbox{log}(\mu)}(\alpha_{SO(16)}(\mu))^{-1}=-\frac{b_{SO(16)}(\mu)}{2\pi}.$ (7.1)
where $b_{SO(16)}(\mu)$ is the $\beta$-function coefficient of $SO(16)$ at the energy scale $\mu$. More explicitly,
$b_{SO(16)}(\mu)=-\frac{11}{3}\sum_{\rm Vector}T_{SO(16)}(R_{V})+\frac{2}{3}\sum_{\rm Weyl}T_{SO(16)}(R_{F})+\frac{1}{6}\sum_{\rm Real}T_{SO(16)}(R_{S}).$ (7.2)
where Vector, Weyl, and Real under the summation symbol stand respectively for real vector, Weyl fermion, and real scalar fields in terms of 4D theories. The summation is done for particles with mass smaller than the specified energy scale $\mu$. The vector bosons are the gauge bosons and thus belong to the adjoint representation 120 of $SO(16)$ with $T(R_{V}={\bf 120})=C_{2}(SO(16))=14$, where $C_{2}(SO(16))$ is the quadratic Casimir invariant of the adjoint representation of $SO(16)$, and $T(R_{X})$ is the Dynkin index of the irreducible representation $R_{X}$ of $SO(16)$. Finally, $T({\bf 128})=16$ [39].
Let us now consider the RGE for the 4D gauge coupling constant in the 6D gauge theories given in Eq. (7.1) by using the $\beta$ function coefficient given in Eq. (7.2), which depends on its field content at the energy scale $\mu$. For the $\beta$ function coefficient, we take into account the contribution from not only the zero modes but also the KK modes with masses less than the renormalization scale $\mu$. Once specifying the mass spectrum of the model, we can calculate the RGE for the gauge coupling constant at one-loop level. In general, it is difficult to write down the exact mass spectrum because it depends on orbifold boundary conditions as well as the parameters of 6D bulk and 4D localized terms. At the leading-order approximation, the masses of zero modes and 4D fields are much less than $\mu$ and those of the $k$-th KK modes can be set as $m=km_{KK}$. By using this mass spectrum, the RGE of the gauge coupling constant can be divided into two regimes:
$\frac{d}{d\mbox{log}(\mu)}\left(\alpha_{SO(16)}(\mu)\right)^{-1}\simeq\left\{\begin{array}[]{ll}-\frac{1}{2\pi}b_{SO(16)}^{0}&\ \ \mbox{for}\ \ \mu<m_{KK}\\ -\frac{1}{2\pi}\left(b_{SO(16)}^{0}+k\Delta b_{SO(16)}^{KK}\right)&\ \ \mbox{for}\ \ km_{KK}\leq\mu<\left(k+1\right)m_{KK}\\ \end{array}\right.,$ (7.5)
where $b_{SO(16)}^{0}$ is the $\beta$-function coefficient derived from the 6D zero modes and 4D fields and $\Delta b_{SO(16)}^{KK}$ is the additional $\beta$-function coefficient induced by the KK modes of all 6D fields, both of which can be calculated by using Eq. (7.2). More explicitly, the latter is given by
$\Delta b_{SO(16)}^{KK}$ $=-\frac{10}{3}C_{2}(SO(16))+\frac{4}{3}\sum_{\rm 6D\ Weyl}T_{SO(16)}(R_{F})$ (7.6)
because each 6D bulk gauge field is decomposed into one 4D gauge boson and two scalars and each 6D bulk Weyl fermion is decomposed into a 4D Dirac fermion. Our model includes one 6D $SO(16)$ gauge boson in the adjoint representation ${\bf 120}$ of $SO(16)$ and two 6D Weyl fermions in the spinor representation ${\bf 128}$ of $SO(16)$. Therefore, Eq. (7.6) gives
$\Delta b_{SO(16)}^{KK}$ $=-\frac{10}{3}\times 14+\frac{4}{3}\times 2\times 16=-4<0.$ (7.7)
Therefore, the $SO(16)$ gauge coupling constant is found to be asymptotically free.
Finally, we comment on the RGE evolutions of the gauge coupling constants at energy scales below the unification scale. At energy scales below the KK scale, the zero modes of the 6D fields and the 4D fields contribute to the RGE evolutions of the gauge coupling constants. As is well known, the contribution of the SM particles alone does not unify the three SM gauge coupling constants at a single high energy scale. Therefore, non-SM particles of a specific mass spectrum or symmetry breaking at scales between the electroweak and unification scales are required to satisfactorily unify the gauge coupling constants. Related discussions can be found in various models [52, 53, 54, 55, 56, 57, 58, 59].
8 Summary and discussions
In this paper, we have proposed a unified model for the SM gauge symmetry and family symmetry based on the $SO(16)$ gauge symmetry on the 6D spacetime $M^{4}\times D^{2}/\mathbb{Z}_{N}$ $(N\geq 9)$. In this model, the three chiral generations of quarks and leptons in the SM can be unified into one 6D Weyl fermion in the spinor representation of $SO(16)$ when considering the symmetry-breaking pattern $SO(16)\to SO(10)\times SU(3)\times U(1)$. The 6D $SO(16)$ gauge anomaly is canceled due to the vectorlike nature of the Weyl fermion in the model, and the 4D gauge anomalies are canceled by introducing suitable 4D localized fermions at the fixed point. The gauge coupling constant of $SO(16)$ is found to be asymptotically free above the KK scale.
We note in passing that an $SO(18)$ model can unify the 6D Weyl fermions into the spinor representation 256 with three chiral generations of the SM fermions without 6D $SO(18)$ gauge anomaly. Using the same arguments for $SO(16)$ given in Sec. 7, the equation corresponding to Eq. (7.6) in such a model is given by
$\Delta b_{SO(18)}^{KK}$ $=-\frac{10}{3}C_{2}(SO(18))+\frac{4}{3}\sum_{\rm 6D\ Weyl}T_{SO(18)}(R_{F})=-\frac{10}{3}\times 16+\frac{4}{3}\times 2\times 32=+32>0,$ (8.1)
where $C_{2}(SO(18))=16$ and $T_{SO(18)}({\bf 256})=32$ [39]. Hence, the gauge coupling constant in the $SO(18)$ model is not asymptotically free. Basically, the cause of this difference arises from the fact that the dimensions of the adjoint and spinor representations increase with increasing $N$ in $SO(N)$.
It is worth mentioning that for gauge-Higgs unification (GHU) models, which unify gauge bosons and Higgs bosons [60, 61, 62, 63, 64], one may use the $SO(N)$ gauge symmetry as in this work to construct a model that embeds $SU(3)_{C}\times SO(5)_{W}\times U(1)$ electroweak GHU models [65, 66, 67] or $SO(11)$ gauge-Higgs GUT models [26, 68, 69, 70, 71, 72]. For phenomenological aspects of the models, see Refs. [73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84].
Finally, we make a brief comment on the masses of quarks and leptons. The analysis of mass matrices in models with $SU(3)$ family symmetry has already been done in Refs. [85, 86, 87, 88, 89, 90, 6, 91, 92, 93, 94, 95, 96]. Though phenomenologically important, a detailed study of masses and mixing in this model, however, is outside the scope of this paper and will be discussed in future work.
Acknowledgments
This work was supported in part by the Japan Society for the Promotion of Science (JSPS) KAKENHI Grant No. JP21H05182 (N.Y.), by the National Science and Technology Council of Taiwan under Grant Nos. NSTC-111-2112-M-002-018-MY3 (C.W.C.), and by the National Key Research and Development Program of China Grant No. 2020YFC2201504, by the Project No. 12275333 supported by the National Natural Science Foundation of China, by the Scientific Instrument Developing Project of the Chinese Academy of Sciences, Grant No. YJKYYQ20190049, and by the International Partnership Program of Chinese Academy of Sciences for Grand Challenges, Grant No. 112311KYSB20210012 (T.L.).